OpenStudy (anonymous):

Solve 12x2 + x - 6 = 0 by factoring

OpenStudy (anonymous):

Notice 4 times 3 equals 12 Notice that 3 times 2 equals 6 We have (4x + 3)(3x - 2) = 0 Now solve 4x + 3 = 0 4x = -3 x = -3/4 and the same for the other one. ..